If (log x)/(2a+3b-5c)=(log y)/(2b+3c-5a)=(log z)/(2c+3a-5b) then xyz=
a.2
b.1
c.0
d.-1

Grade:12th pass

1 Answers

Ram Gudikandula
6 years ago
Given:- logx/2a+3b-5c=logy/2b+3c-5a=logx/2c+3a-5bTherefore, we can write it as logx+logy+ logz/0So, logx+ logy+ logz=0So,log(xyz)=0xyz=1